Find the complex zeros of the polynomilal function:
f(x)=x^4-8x^3+16x^2+8x-17

Answer by stanbon(48535)

f(x)=x^4-8x^3+16x^2+8x-17
Since the coefficients add to zero, x=1 is a zero.
Use synthetic division to find the other factor which
is x^3-7x^2+9x+17
x=-1 is a zero of this cubic leaving a factor of
x^2-8x+17
Then, using the quadratic formula you get:
x=[8+-sqrt(8^2-4*17)]/2
x=[8+-sqrt(-4)]/2
x=[-4+-i]
Zeroes are : x=1, x=-1, x=-4+i, x=-4-i
Cheers,
Stan H.